# Integrating Intent Understanding and Optimal Behavior Planning for Behavior Tree Generation from Human Instructions

### v1

@@ 1 Introduction
Behavior Trees (BTs) employ a logical tree structure to control various high-level robot behaviors, which offer several advantages, including modularity, interpretability, and reactivity. Due to these characteristics, BTs have become a popular robot control architecture in recent years. Moreover, approaches to automatically generate BTs have also been widely studied, including evolutionary computing Colledanchise et al. (2019b), reinforcement learning Banerjee (2018), and BT synthesis Li et al. (2021). BT Expansion Cai et al. (2021) is a sound and complete algorithm for constructing BTs through behavior planning, theoretically ensuring the success of the BT. However, existing methods typically generate BTs from fitness functions, reward functions, or formalized goal conditions, but few studies have focused on generating BTs directly from natural language instructions.
@@
Our experiments demonstrate that the LLM (ERNIE 4.0), powered by few-shot demonstrations and reflective feedback, achieves high performance in both grammar and interpretation accuracy for intent understanding. The analysis of optimal behavior planning shows that the OBTEA produces BTs with lower costs and fewer condition node ticks compared to the baseline BT Expansion algorithm. Finally, the success of deployment in a café scenario highlights the promising future of our framework.
@@ 3.3 Optimal Behavior Planning
Table 1: Performance analysis of ERNIE 4.0 on intent understanting in the café sencario.
Difficulty Instructions Examples Goals Demonstrations GA-NF(%) GA-1F(%) GA-5F(%) IA(%)
Easy 30 Please come to the bar. RobotNear(Bar) Zero-Shot 51.7 54.9 64.9 26.7
One-Shot 88.3 93.3 98.3 83.3
Five-Shot 92.7 99.9 99.9 91.9
Medium 30 Please make sure there’s enough water and the floor stays clean. Exist(Water) $\wedge$ $\neg$ Dirty(Floor) Zero-Shot 49.1 54.3 71.3 20.4
One-Shot 81.1 88.9 98.9 75.6
Five-Shot 91.1 99.9 99.9 83.3
Hard 40 I’m at table three, bring some dessert and yogurt. If there is no yogurt, coffee is fine. On(Dessert,Table3) $\wedge$ (On(Yogurt,Table3) $\vee$ On(Coffee,Table3)) Zero-Shot 38.3 52.5 65.0 12.5
One-Shot 72.5 84.9 88.3 50.8
Five-Shot 87.1 92.3 94.9 76.9
@@ 4.3 Analysis for Optimal Behavior Planning
Table 3 shows the results averaged over 1000 runs. The results again emphasize OBTEA’s proficiency in achieving lower action costs and condition node ticks across all cases with varying complexity. In particular, in the most complex case 9, OBTEA significantly reduces costs by more than $50\%$, and condition node ticks decrease by approximately $30\%$. When there are more available paths to reach the goal, the costs of BTs found by OBTEA are smaller than the baseline. It is worth noting that while OBTEA has achieved improvements, its additional time overhead is also acceptable. The test results and discussions regarding planning time are in the appendix.
@@ 5 Related Work
The study of automatic BT generation is a significant area of research in robotics and AI. Various methods have been explored for this purpose. Heuristic search methods, such as grammatical evolution Neupane and Goodrich (2019), genetic programming Colledanchise et al. (2019b); Lim et al. (2010), and Monte Carlo DAG Search Scheide et al. (2021), have been widely investigated. In addition to heuristic approaches, machine learning techniques have been applied to generate BTs. In particular, methods such as reinforcement learning Banerjee (2018); Pereira and Engel (2015) and imitation learning French et al. (2019) have gained attention. There are also research efforts that employ formal synthesis methods to generate BTs, such as Linear Temporal Logic (LTL) Li et al. (2021) and its variant approaches Tadewos et al. (2022); Neupane and Goodrich (2023). However, it is worth noting that this approach can be resource intensive in terms of modeling the environment and the computational resources required to solve the problem. Last but not least, there are planning-based methods. Approaches based on STRIPS-style modeling Fikes and Nilsson (1971) offer more intuitive environment modeling and can ensure the rationality of behavior tree structures Colledanchise et al. (2019a); Styrud et al. (2022). In Cai et al. (2021), a formalism based on the STRIPS style is introduced, called BT expansion, which is a complete and sound algorithm for behavior planning.

### v2

@@ 1 Introduction
Behavior Trees (BTs) employ a logical tree structure to control various high-level robot behaviors, which offer many advantages, including modularity, interpretability, reusability, reactivity, and robustness. Due to these characteristics, BTs have become a popular robot control architecture in recent years. Moreover, approaches to automatically generate BTs have also been widely studied, including evolutionary computing Colledanchise et al. (2019), reinforcement learning Banerjee (2018), and BT synthesis Li et al. (2021). BT Expansion Cai et al. (2021) is a sound and complete algorithm for constructing BTs through behavior planning, theoretically ensuring the success of the BT. However, existing methods typically generate BTs from fitness functions, reward functions, or formalized goal conditions, but few studies have focused on generating BTs directly from natural language instructions.
@@
Our experiments demonstrate that the LLM (GPT 3.5), powered by few-shot demonstrations and reflective feedback, achieves high performance in both grammar and interpretation accuracy for intent understanding. The analysis of optimal behavior planning shows that the OBTEA produces BTs with lower costs and fewer condition node ticks compared to the baseline BT Expansion algorithm. Finally, the success of deployment in a café scenario highlights the promising future of our framework.
@@ 3.3 Optimal Behavior Planning
Difficulty Data Examples Goals Demonstrations GA-NF(%) GA-1F(%) GA-5F(%) IA(%)
Easy 30 Please come to the bar. RobotNear(Bar) Zero-Shot 60.0 74.7 91.3 56.7
Few-shot-1 87.3 95.3 100.0 85.3
Few-shot-5 92.0 96.7 100.0 90.7
Medium 30 Ensure there’s enough water and keep the hall light on. Exist(Water) $\wedge$ Active(HallLight) Zero-Shot 50.0 66.0 88.0 57.3
Few-shot-1 77.3 91.3 98.7 62.0
Few-shot-5 92.7 94.7 99.3 82.0
Hard 40 I’m at table three, please bring yogurt and keep the AC warm. If there is no yogurt, coffee is fine. (On(Yogurt,Table3) $\wedge$ (On(Coffee,Table3)) $\vee$ $\neg$ Low(ACTemperature) Zero-Shot 48.0 62.5 84.0 42.0
Few-shot-1 76.5 87.5 93.0 51.0
Few-shot-5 80.5 88.0 96.0 65.0
Table 1: Performance analysis of GPT 3.5 on intent understanting in the café sencario.
@@ 4.3 Analysis for Optimal Behavior Planning
Table 3 shows the results averaged over 1000 runs. The results again emphasize OBTEA’s proficiency in achieving lower action costs and condition node ticks across all cases with varying complexity. In particular, in the most complex case 9, OBTEA significantly reduces costs by more than $50\%$, and condition node ticks decrease by approximately $30\%$. When there are more available paths to reach the goal, the costs of BTs found by OBTEA are smaller than the baseline. Note that while OBTEA has achieved improvements, its additional time overhead is also acceptable. The test results and discussions regarding planning time are in the appendix.
@@ 5 Related Work
The study of automatic BT generation is a significant area of research in robotics and AI. Various methods have been explored for this purpose. Heuristic search methods, such as grammatical evolution Neupane and Goodrich (2019), genetic programming Colledanchise et al. (2019); Lim et al. (2010), and Monte Carlo DAG Search Scheide et al. (2021), have been widely investigated. In addition to heuristic approaches, machine learning techniques have been applied to generate BTs. In particular, methods such as reinforcement learning Banerjee (2018); Pereira and Engel (2015) and imitation learning French et al. (2019) have gained attention. There are also research efforts that employ formal synthesis methods to generate BTs, such as Linear Temporal Logic (LTL) Li et al. (2021) and its variant approaches Tadewos et al. (2022); Neupane and Goodrich (2023). However, it is worth noting that this approach can be resource intensive in terms of modeling the environment and the computational resources required to solve the problem.
@@ Acknowledgments
Contribution Statement
The contributions of Xinglin Chen and Yishuai Cai to this paper were equal. Corresponding author: Minglong Li.